# Distributed Differentially Private Data Analytics via Secure Sketching

### v1

@@ Abstract.
We explore the use of distributed differentially-private computations across multiple servers, balancing the tradeoff between the error introduced by the differentially-private mechanism and the computational efficiency of the resulting distributed algorithm.
We introduce the linear-transformation model, where clients have access to a trusted platform capable of applying a public matrix to their inputs. Such computations can be securely distributed across multiple servers using simple and efficient secure multiparty computation techniques.
The linear-transformation model serves as an intermediate model between the highly expressive central model and the minimal local model. In the central model, clients have access to a trusted platform capable of applying any function to their inputs. However, this expressiveness comes at a cost, as it is often expensive to distribute such computations, leading to the central model typically being implemented by a single trusted server. In contrast, the local model assumes no trusted platform, which forces clients to add significant noise to their data. The linear-transformation model avoids the single point of failure for privacy present in the central model, while also mitigating the high noise required in the local model.
We demonstrate that linear transformations are very useful for differential privacy, allowing for the computation of linear sketches of input data. These sketches largely preserve utility for tasks such as private low-rank approximation and private ridge regression, while introducing only minimal error, critically independent of the number of clients. Previously, such accuracy had only been achieved in the more expressive central model.
@@ 1. Introduction
Differential Privacy (DP) (DMNS, 06) has become the de-facto standard for ensuring the privacy of individuals whose data is used in data analytics. DP offers strong, provable guarantees, such as composability and resilience to auxiliary information. In the classic central model of differential privacy, clients submit their data to a trusted central server, which processes the data and releases a (necessarily) noisy version of the result. Analysts can then use this result to perform queries on the data. Informally, differential privacy guarantees that the analyst cannot confidently determine whether any individual contributed their data, except with some small probability.
@@
Another option is to avoid relying on external servers altogether. In this case, each client must publish their data in a locally differentially private manner, by adding noise directly to their data. While this removes the need for a trusted server, it results in lower utility due to the increased noise. This approach is known as the local model of differential privacy (KLN+, 08), an example of which is randomized response (War, 65).
@@
A notable example of such restricted expressiveness is the class of functions known as shuffles, where a central entity (or a distributed system) randomly permutes the messages from clients before they reach the analysing server. This is known as the shuffle model (BEM+, 17; EFM+, 19; CSU+, 19). The shuffle model is motivated by cryptographic protocols (e.g., mix-nets), which allow for securely distributing the shuffle across multiple servers. The shuffle model has been used to implement many differentially-private mechanisms with reasonable utility (EFM+, 19; CSU+, 19). Surprisingly, the shuffle model is more expressive than initially expected: with a trusted shuffler, any function can be securely computed in two rounds under the assumption of an honest-majority (BHNS, 20).
@@ 1.1. Our Contribution
In this work, we investigate the power of the linear-transformation model (LTM) for differentially-private mechanisms. In this model, the clients only have access to a trusted platform for performing arbitrary linear transformations of their inputs. Linear functions can be distributed extremely efficiently with secure computations. However, it is also known that the expressiveness of this class of functions is strictly limited.
For a visualization of the architecture, we refer to Figure 1. We demonstrate the benefits of this linear transformation model (LTM) by showing that it can be used, in combinations with linear sketches to construct private summaries of a dataset with noise comparable to that of central differentially private mechanisms.
As a warm-up, we show how to use linear sketches to compute vector statistics as described in Section 3.1. For more advanced application, we consider problems from linear algebra such as low rank approximation and regularized linear regression. Note that in the LTM no noise is added centrally, and therefore clients (like in the local model) must add some noise to their data before sending it to the trusted platform. However, we show that thanks to the centrally applied linear transformation, we can significantly reduce the noise to be added (and therefore obtain better utility). Although it is not inherent to the model, all of our mechanisms require only a single round of communication from clients to server, akin to streaming algorithms.
In a nutshell, we obtain utility guarantees similar to the central model, while using significantly weaker assumptions that the central or shuffle model. A full overview of related work is given in Table 1.
Low Rank Approximation.
We are given a data matrix $\mathbf{A}\in\mathbb{R}^{n\times d}$. Our goal is to compute an orthogonal projection $\mathbf{X}\in\mathbb{R}^{d\times r}$ minimizing the error $OPT=\min_{\mathbf{X}}\|\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}^{2}$, where .$\|.\|_{F}$ denotes the Frobenius norm of a matrix. In the LTM, we can compute an $\hat{\mathbf{X}}$ differentially privately with error $\|A-\hat{X}\hat{X}^{T}A\|_{F}^{2}\leq(1+o(1))\cdot(\|A-XX^{T}A\|_{F}^{2}+\text{poly}(\varepsilon^{-1},d,k,\log 1/\delta)$. Notably, this bound is independent of $n$, which so far has only been achieved in the more expressive central model.
Ridge Regression.
@@
Comparison to state-of-the-art.
Table 1 provides a comparison of the utility bounds for the problems in the local, central, shuffle and LTM model. We obtain similar utility to mechanisms run in the more expressive central model, while at the same time we are able to distribute the evaluation of the mechanism and avoid the single point of failure of the central model. When comparing with the similar shuffle model for frequency estimation, it should be noted that the LTM is a strictly less expressive model. Indeed, it is possible to use the shuffle model to emulate the LTM, while the converse is not possible (remember that the matrix is public in the LTM, while the permutation is secret in the shuffle model). For low rank approximation/PCA, we are not aware of any work in the shuffle model. All of our mechanisms require only a single round of communication between users and servers, with one messages per user sent to each server in that round.
1.2. Related Work
@@
There are various ways in which one could formulate the low rank approximation problem. The setting which is most important to us is the seminal paper (DNMR, 14), who achieved a worst case additive error of the order $d\sqrt{k}$ for outputting an orthogonal projection matrix $V_{k}$ in the row space of $\mathbf{A}$. This bound is also optimal. In the local model where each client holds a row of the data matrix $\mathbf{A}$, (Upa, 18) gave an algorithm with an additive error of the order $\sqrt{n}$, which matches the lower bound by (BS, 15) up to lower order terms. To the best of our knowledge, no single round shuffle protocol improving over the local bounds is known.
Low rank approximation and its sister problem PCA, have seen substantial attention in settings more loosely related to our work. Much of the work on private low rank approximation (BDMN, 05; CSS, 12; DTTZ, 14; HR, 12, 13; HP, 14; BDWY, 16; KT, ) considers the data to be fixed, often using Gaussian noise. In addition, a strong spectral gap assumption, typically of the form that the first singular value is substantially larger than the second, is crucial to the analysis. We note that all algorithms operating with this assumption do not yield worst case bounds.
When there is no spectral gap, we compare to the central model exponential mechanism approach from (CSS, 12), implemented by (KT, ). (LKO, 22) also provide a solution to this problem; however, it is unfortunately computationally intractable.
@@
The previous work most related to ours is due to (She, 19). The author gave a mechanism that preserved the entire spectrum of a data matrix in a private manner and also showed that the returned regression vector $\hat{\mathbf{x}}$ is close to the optimal $\mathbf{x}$ assuming that the matrix is well conditioned.
Most previous work (KST, 12; CWZ, 21; BST, 14; WFS, 15; WGX, 18; WSX, 18; WHZ+, 23; Wan, 18; VTJ, 22; MKFI, 22; LKO, 22; ZMW, 17) study private linear regression in the context of risk minimization, where the algorithm is given i.i.d. samples from some unknown distribution and aims at computing a solution with good out of sample performance. Even without privacy constraints risk minimization yields an additive error that depends on $n$. Another line of work studies varies loss functions, including regression in a Bayesian setting (MASN, 16; FGWC, 16; DNMR, 14). We are not aware of any prior work on linear regression in the local model of differential privacy.
Differential Privacy via Johnson-Lindenstrauss Transforms.
The Johnson-Lindenstrauss lemma has been previously used in differential privacy. (BBDS, 12) studies the DP guarantee a JL transform itself gives in the central model, in the context of cut queries and directional variance queries. (KKMM, 13) and (Sta21a, ) use JL transforms together with Gaussian noise in the context of differentially private Euclidean distance approximation, to decrease the error. This was subsequently improved in (Sta21b, ), using Laplace noise. The difference to our work is that the transform is used to reduce $d$ instead of $n$, by having every client apply it locally before adding Gaussian noise. In (Nik, 23) the authors make use of JL transforms to achieve private query release, by applying it before releasing a bundle of queries. (GKK+, 23) studies pairwise statistics in the local model of differential privacy and uses JL transforms to reduce client-sided dimensionality.
Local Shuffle LTM (this work) Central
Frequency Est $\tilde{O}(\sqrt{n})$ (BNST, 20) $\tilde{\Theta}(\varepsilon^{-1})$(GGK+, 21) $\tilde{O}(\varepsilon^{-2})$ $\tilde{\Theta}(\varepsilon^{-1})$
$\Omega(\sqrt{n})$ (BS, 15)
Low-Rank Approximation $\tilde{O}(\sqrt{n})$(Upa, 18) N.A. $\tilde{O}(d^{3}k^{2}\varepsilon^{-2}\log\frac{d}{\delta})$ $\tilde{\Theta}(d\sqrt{k}\varepsilon^{-2})$(DTTZ, 14)
PCA $\Omega(\sqrt{n})$ (BS, 15)
Table 1. SOTA utility bounds for various models for frequency estimation and low rank approximation.
@@ 2. Preliminaries
Column vectors are written in bold lowercase letters $\mathbf{b}$ and matrices in bold uppercase letters $\mathbf{A}$. The transpose operator over vectors and matrices is $\mathbf{b}^{T}$ and $\mathbf{A}^{T}$. For any vector $\mathbf{b}$, we denote by $\lVert\mathbf{b}\rVert=\sqrt{\sum_{i}b_{i}^{2}}$ its $\ell_{2}$-norm. For any matrix $\mathbf{A}$, we denote by $\lVert\mathbf{A}\rVert_{F}=\sqrt{\sum_{i}\sum_{j}\mathbf{A}_{i,j}^{2}}$ its Frobenius norm. We denote the inner product between vectors $\langle\mathbf{a},\mathbf{b}\rangle=\mathbf{a}^{T}\mathbf{b}$. Two vectors $\mathbf{a},\mathbf{b}$ are orthogonal if their inner product is 0. An orthogonal matrix $\mathbf{A}$ is a real matrix where the columns have unit Euclidean norm and are pairwise orthogonal.
Subspace Preserving Sketches.
@@
We consider subspace-preserving sketches in this work. Of particular interest for us are algorithms that compute a subspace approximation of $A$ without prior knowledge of $A$. Such algorithms are known as oblivious subspace embeddings. We will use a family of such embeddings known as oblivious sparse norm-approximating projections originally due to (NN, 13), but since improved upon in several subsequent works.
Definition 2.1 (OSNAP (NN, 13).).
@@
Let $\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ be a distribution over random matrices $\{-1,0,1\}^{m\times n}$, where $s$ entries of each column, chosen uniformly and independently for each column, are set to ±$\pm 1$ with equal probability, and all other entries are set to 0. If $s=1$, we write $\mathcal{D}_{m,n}^{\mathtt{sketch}}$. We say that $S\sim\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ is an $(\alpha,\beta,s,m)$ OSNAP for a $d$-dimensional subspace $W\in\mathbb{R}^{n}$ if with probability $1-\beta$, for all $a\in W$ and some precision parameter $\alpha$
$\left|\|Sa\|_{2}^{2}-\|a\|_{2}^{2}\right|\leq\alpha\cdot\|a\|^{2}.$
@@
We give bounds on available choices of $(\alpha,\beta,s,m)$ for subspaces of rank $k$ in table 5 in the appendix.
Differential Privacy.
@@
Differential privacy (DMNS, 06) offers privacy guarantees to individuals contributing their data to some randomized algorithm. We say that two datasets are neighboring if one can be obtained from the other by the the replacement of a single individual with another individual.
Definition 2.2 (Differential Privacy in the central model (DMNS, 06)).
@@
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\varepsilon,\delta)$ differentially private, if for all neighboring data sets $x,x^{\prime}\in\mathcal{X}$ and all outputs $S\subseteq\mathcal{Y}$ it holds that
@@
where the probabilities are over the randomness of $\mathcal{M}$.
Lemma 2.3 (The Gaussian Mechanism (DR, 14)).

### v2

@@ Abstract.
We introduce the linear-transformation model, a distributed model of differentially private data analysis. Clients have access to a trusted platform capable of applying a public matrix to their inputs. Such computations can be securely distributed across multiple servers using simple and efficient secure multiparty computation techniques.
The linear-transformation model serves as an intermediate model between the highly expressive central model and the minimal local model. In the central model, clients have access to a trusted platform capable of applying any function to their inputs. However, this expressiveness comes at a cost, as it is often prohibitively expensive to distribute such computations, leading to the central model typically being implemented by a single trusted server. In contrast, the local model assumes no trusted platform, which forces clients to add significant noise to their data. The linear-transformation model avoids the single point of failure for privacy present in the central model, while also mitigating the high noise required in the local model.
We demonstrate that linear transformations are very useful for differential privacy, allowing for the computation of linear sketches of input data. These sketches largely preserve utility for tasks such as private low-rank approximation and private ridge regression, while introducing only minimal error, critically independent of the number of clients.
@@ 1. Introduction
Differential Privacy (DP) (Dwork et al., 2006b) has become the de-facto standard for ensuring the privacy of individuals whose data is used in data analytics. DP offers strong, provable guarantees, such as composability and resilience to auxiliary information. In the classic central model of differential privacy, clients submit their data to a trusted central server, which processes the data and releases a (necessarily) noisy version of the result. Analysts can then use this result to perform queries on the data. Informally, differential privacy guarantees that the analyst cannot confidently determine whether any individual contributed their data, except with some small probability.
@@
Another option is to avoid relying on external servers altogether. In this case, each client must publish their data in a locally differentially private manner, by adding noise directly to their data. While this removes the need for a trusted server, it results in lower utility due to the increased noise. This approach is known as the local model of differential privacy (Kasiviswanathan et al., 2008), an example of which is randomized response (Warner, 1965).
@@
A notable example of such restricted expressiveness is the class of functions known as shuffles, where a central entity (or a distributed system) randomly permutes the messages from clients before they reach the analysing server. This is known as the shuffle model (Bittau et al., 2017; Erlingsson et al., 2019; Cheu et al., 2019). The shuffle model has been used to implement many differentially-private mechanisms with reasonable utility (Erlingsson et al., 2019; Cheu et al., 2019). Surprisingly, the shuffle model is more expressive than initially expected: with a trusted shuffler, any function can be securely computed in two rounds under the assumption of an honest-majority (Beimel et al., 2020).
@@ 1.1. Our Contribution
In this work, we investigate the power of the linear-transformation model (LTM) for differentially-private mechanisms. In this model, the clients only have access to a trusted platform for performing arbitrary linear transformations of their collective inputs. Linear functions can be distributed extremely efficiently with secure computations. However, it is also known that the expressiveness of this class of functions is strictly limited.
For a visualization of the architecture, we refer to Figure 1. We demonstrate the benefits of this linear transformation model (LTM) by showing that it can be used, in combinations with linear sketches to construct private summaries of a data set with noise comparable to that of central differentially private mechanisms. As in the local model, each client adds some amount of noise to their data before passing it onto the linear sketch. We then release the linear transform of the data. The benefit of having a trusted linear sketch is that the noise can be combined, thereby significantly reducing the amount of noise per client. All of our mechanisms require only a single round of communication from clients to server.
We demonstrate the effectiveness of this model by exemplarily applying it to problems in numerical linear algebra. A particularly powerful and expressive class of linear sketches with numerous linear algebra applications (Woodruff, 2014) are Johnson-Lindenstrauss (JL) transforms. In this paper, we focus on low rank approximation and regularized linear regression. Our contribution is to show how to design mechanisms for these problems, with the most widely used JL-transforms as the underlying secure sketch. For these problems, we achieve utility comparable to that of the central model with local privacy guarantees, assuming the JL-transform is computed securely.
For both of the following problems, we have the option of achieving $\varepsilon$ differential privacy, via protocols based on the Laplace mechanism, as well as $(\varepsilon,\delta)$-differential privacy with slightly higher utility, via protocols based on the Gaussian mechanism.
Low Rank Approximation. We are given a data matrix $\mathbf{A}\in\mathbb{R}^{n\times d}$. Our goal is to compute an orthogonal projection $\mathbf{X}\in\mathbb{R}^{d\times r}$ minimizing the error $OPT=\min_{\mathbf{X}}\|\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}$, where .$\|.\|_{F}$ denotes the Frobenius norm of a matrix. In the LTM, we give a $(\varepsilon,\delta)$ differentially private mechanism with multiplicative error $(1+\alpha)$ and additive error $\tilde{O}(kd^{3/2}\alpha^{-3}\varepsilon^{-1}\log\delta^{-1})$. We also present a $\varepsilon$ differentially private mechanism with multiplicative error $(1+\alpha)$ and additive error $\tilde{O}(k^{3}d^{3/2}\alpha^{-6}\varepsilon^{-1})$. State of the art central model mechanisms (Dwork et al., 2014; Hardt and Roth, 2012) have no multiplicative error and additive error $O(k\sqrt{d}\cdot\varepsilon^{-1})$ , which is optimal in terms of $k$ and $d$. For the local model, we are only aware of the result by (Arora et al., 2018), which has a multiplicative error of $O(\log n\log^{3}d)$ and additive error $O(k^{2}n\cdot d\text{poly}(\varepsilon^{-1},\log\delta^{-1}))$. There are several related questions we will survey in more detail in Section 2.
@@
Table 1. SOTA utility bounds for various models for frequency estimation and low rank approximation. Our bounds for low rank approximation have a multiplicative error $O(1+\alpha_{\mathbf{S}})$, which we omitted for space.
Local Shuffle LTM (this work) Central
Frequency Est $\tilde{O}(\sqrt{n})$(Bassily et al., 2020) $\tilde{\Theta}(\varepsilon^{-1})$ (Ghazi et al., 2021) $\tilde{O}(\varepsilon^{-1})$ $\tilde{\Theta}(\varepsilon^{-1})$
$\Omega(\sqrt{n})$(Bassily and Smith, 2015)
Low-Rank App $(\tilde{O}(\log n\log^{3}d),O(k^{2}nd))$(Arora et al., 2018) N.A. $\tilde{O}(\frac{kd^{3/2}}{\alpha_{\mathbf{S}}^{3}}\varepsilon^{-1})$ $\tilde{\Theta}(k\sqrt{d}\varepsilon^{-1})$(Dwork et al., 2014)
$\Omega(\sqrt{n})$(Bassily and Smith, 2015)
Table 1 provides a comparison of the utility bounds for the problems in the local, central, shuffle and LTM models. To the best of our knowledge, there are no results on low rank approximation in the shuffle model.
As a warm-up to illustrate the potential inherent in the LTM, and to also highlight the comparison to the other models, we consider the basic problem of computing histograms and frequency of moments.
2. Related Work
@@
There are various ways in which one could formulate the low rank approximation problem. The setting which is most important to us is the seminal paper (Dwork et al., 2014), who achieved a worst case additive error of the order $k\sqrt{d}$ for outputting an orthogonal projection matrix $V_{k}$ in the row space of $\mathbf{A}$. This bound is also optimal. In the local model where each client holds a row of the data matrix $\mathbf{A}$, and we wish to output an orthogonal projection in the column space of $\mathbf{A}$, (Upadhyay, 2018) gave an algorithm with an additive error of the order $\sqrt{n}$, which matches the lower bound by (Bassily and Smith, 2015) up to lower order terms. This is different from computing a projections in the row space of $\mathbf{A}$ considered by (Dwork et al., 2014) and indeed our own results. The only know local algorithm to do likewise in the local model is due to (Arora et al., 2018), with a substantially larger multiplicative and additive error compared to our results. To the best of our knowledge, no single round shuffle protocol improving over the local bounds is known.
Low rank approximation and its sister problem PCA, have seen substantial attention in settings more loosely related to our work. Much of the work on private low rank approximation (Blum et al., 2005; Chaudhuri et al., 2012; Dwork et al., 2014; Hardt and Roth, 2012, 2013; Hardt and Price, 2014; Balcan et al., 2016; Kapralov and Talwar, [n. d.]) considers the data to be fixed, often using Gaussian noise. In addition, a strong spectral gap assumption, typically of the form that the first singular value is substantially larger than the second, is crucial to the analysis. We note that all algorithms operating with this assumption do not yield worst case bounds.
When there is no spectral gap, we compare to the central model exponential mechanism approach from (Chaudhuri et al., 2012), implemented by (Kapralov and Talwar, [n. d.]). (Liu et al., 2022) also provide a solution to this problem; however, it is unfortunately computationally intractable.
@@
The previous work most related to ours is due to (Sheffet, 2019). The author gave a mechanism that preserved the entire spectrum of a data matrix in a private manner and also showed that the returned regression vector $\hat{\mathbf{x}}$ is close to the optimal $\mathbf{x}$ assuming that the matrix is well conditioned.
Most previous work (Kifer et al., 2012; Cai et al., 2021; Bassily et al., 2014; Wang et al., 2015, 2018a, 2018b, 2023; Wang, 2018; Varshney et al., 2022; Milionis et al., 2022; Liu et al., 2022; Zheng et al., 2017) study private linear regression in the context of risk minimization, where the algorithm is given i.i.d. samples from some unknown distribution and aims at computing a solution with good out of sample performance. Even without privacy constraints risk minimization yields an additive error that depends on $n$. Another line of work studies varies loss functions, including regression in a Bayesian setting (Minami et al., 2016; Foulds et al., 2016; Dimitrakakis et al., 2014). We are not aware of any prior work on linear regression in the local DP model.
DP via Johnson-Lindenstrauss Transforms.
The Johnson-Lindenstrauss lemma has been previously used in differential privacy. (Blocki et al., 2012) studies the DP guarantee a JL transform itself gives in the central model, in the context of cut queries and directional variance queries. (Kenthapadi et al., 2013) and (Stausholm, 2021a) use JL transforms together with Gaussian noise in the context of DP Euclidean distance approximation, to decrease the error. This was subsequently improved in (Stausholm, 2021b), using Laplace noise. The difference to our work is that the transform is used to reduce $d$ instead of $n$, by having every client apply it locally before adding noise. In (Nikolov, 2023) the authors make use of JL transforms to achieve private query release, by applying it before releasing a bundle of queries. (Ghazi et al., 2023) studies pairwise statistics in the local model of differential privacy and uses JL transforms to reduce client-sided dimensionality.
Differentially Private Non-Oblivious Sketches.
Non-Oblivious sketches, such as CountSketch, have also been used in combination with either distributed or central DP. (Zhao et al., 2022; Pagh and Thorup, 2022) provide an analysis of differentially private CountSketch in the central model. (Melis et al., 2016) consider differentially private CountSketch and Count-Min Sketch in a distributed setting, but they reduce the dimensionality of each input, which can be done by users locally, rather than reducing the dependence on the number of users, as we do.
@@ 3. Preliminaries
Notation. Column vectors are written in bold lowercase letters $\mathbf{b}$ and matrices in bold uppercase letters $\mathbf{A}$. The transpose operator over vectors and matrices is $\mathbf{b}^{T}$ and $\mathbf{A}^{T}$. For any vector $\mathbf{b}$, we denote by $\lVert\mathbf{b}\rVert=\sqrt{\sum_{i}b_{i}^{2}}$ its $\ell_{2}$-norm. For any matrix $\mathbf{A}$, we denote by $\lVert\mathbf{A}\rVert_{F}=\sqrt{\sum_{i}\sum_{j}\mathbf{A}_{i,j}^{2}}$ its Frobenius norm. We denote the inner product between vectors $\langle\mathbf{a},\mathbf{b}\rangle=\mathbf{a}^{T}\mathbf{b}$. Two vectors $\mathbf{a},\mathbf{b}$ are orthogonal if their inner product is 0. An orthogonal matrix $\mathbf{A}$ is a real square matrix where the columns have unit Euclidean norm and are pairwise orthogonal.
@@
Subspace Preserving Sketches. We consider subspace-preserving sketches in this work, which are sparse Johnson-Lindenstrauss transforms. Of particular interest for us are algorithms that compute a subspace approximation of $A$ without prior knowledge of $A$. Such algorithms are known as oblivious subspace embeddings. We will use a family of such embeddings known as oblivious sparse norm-approximating projections originally due to (Nelson and Nguyên, 2013), but since improved upon in several subsequent works.
Definition 3.1 (OSNAP (Nelson and Nguyên, 2013).).
@@
Let $\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ be a distribution over random matrices $\{-1,0,1\}^{m\times n}$, where $s$ entries of each column, chosen uniformly and independently for each column, are set to ±$\pm 1$ with equal probability, and all other entries are set to 0. If $s=1$, we write $\mathcal{D}_{m,n}^{\mathtt{sketch}}$. We say that $S\sim\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ is an $(\alpha,\beta,s,m)$ OSNAP for a $n\times d$ orthogonal matrix $\mathbf{W}$ if with probability $1-\beta$, for all $a,b\in\mathbb{R}^{d}$ and some precision parameter $\alpha$
$\left|a^{T}\mathbf{W}^{T}\mathbf{S}^{T}\mathbf{S}\mathbf{W}b-a^{T}\mathbf{W}^{T}\mathbf{W}b\right|\leq\alpha\cdot\|\mathbf{W}a\|\|\mathbf{W}b\|.$
@@
We give bounds on available choices of $(\alpha,\beta,s,m)$ for subspaces of rank $k$ in Table 2.
Table 2. Trade-off between the target dimension $m$ and sparsity $s$ of sparse OSEs that are generated such that they have $s$ non-zeros entries per column. Here $\alpha$ denotes the accuracy of the OSE and $\beta$ denotes its fail probability. Variable $k$ is the parameter to $k$-rank approximation, which can be replaced by dimension $d$ for linear regression. The paper by (Achlioptas, 2003) was the first to propose the construction, while its application to preserving subspaces was given by (Clarkson and Woodruff, 2009).
Paper Target Dimension $m$ Sparsity $s$ With Probability
(Achlioptas, 2003) $O(\frac{k+\log 1/\beta}{\alpha^{2}})$ $m$ $1-\beta$
(Nelson and Nguyên, 2013) $O(\frac{k^{2}}{\alpha^{2}\beta})$ 1 $1-\beta$
(Cohen, 2016) $O(\frac{k\log(k/\beta)}{\alpha^{2}})$ $O(\frac{\log(k/\beta)}{\alpha})$ $1-\beta$
(Høgsgaard et al., 2023) $O(\frac{k}{\alpha^{2}})$ $O(\frac{1}{\alpha}\cdot(\frac{k}{\log(1/\alpha)}+k^{2/3}\log^{1/3}(k)))$ $1-2^{-k^{2/3}}$
Oblivious subspace embeddings (OSEs) (Sarlós, 2006), allow for faster approximation algorithms for problems in linear algebra. Achlioptas (Achlioptas, 2003) provided the first Johnson-Lindenstrauss transform with some amount of sparsity. The first embedding with asymptotically smaller number of non-zeros than dense Johnson-Lindenstrauss transforms was probably due to (Kane and Nelson, 2012), who also applied them in the context of subspace embeddings. In a remarkable result, the super-sparse version with only $1$ non-zero entries per column, was first analyzed in (Clarkson and Woodruff, 2013) and improved independently in (Meng and Mahoney, 2013; Nelson and Nguyên, 2013), by increasing the target dimension $m$. In (Nelson and Nguyên, 2013), the sparse Johnson-Lindenstrauss transform (Kane and Nelson, 2012) was studied with a wider range of parameters. Subsequent works (Cohen, 2016; Høgsgaard et al., 2023) further analyze and improve the relationship between $m$ and $s$, emphasizing different parameters. Table 2 provides the exact interplay between $m$ and $s$ in those works. (Liang et al., 2014) proposes an approach to boost the success probability $\beta$ of an OSE, which gives an alternative to increasing the target dimension.
When dealing with rank $k$ approximation, we will condition on
$(1-\alpha_{S})\|\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}^{2}$ $\leq\|\mathbf{S}(\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T})\|_{F}^{2}$
$\leq(1+\alpha_{S})\|\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}^{2}$
$(1-\alpha_{S})\|\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\|^{2}$ $\leq\|\mathbf{S}(\mathbf{A}\mathbf{x}-\mathbf{b})\|^{2}$
$\leq(1+\alpha_{S})\|\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\|^{2}$
We further will give our utility proofs using the sparsity/target dimension bounds from (Cohen, 2016). Other tradeoffs are possible, but they have worse bounds for most ranges of parameters. The results in (Meng and Mahoney, 2013; Nelson and Nguyên, 2013; Høgsgaard et al., 2023) give different trade-offs depending on which set of parameters are considered the most dominant. Notably, the $s=1$ sketches of (Meng and Mahoney, 2013; Nelson and Nguyên, 2013) result in an additive error of $\tilde{O}\left(\frac{k^{4}d^{3}}{\alpha_{S}^{4}\beta^{2}}\varepsilon^{-2}\log\frac{1}{\delta}\right)$ for low rank approximation and a term $\tilde{O}\left(\frac{d^{7}\varepsilon^{-2}\log\frac{1}{\delta}}{\alpha_{S}^{5}\beta^{2}\lambda}+\frac{d^{14}\varepsilon^{-4}\log^{2}\frac{1}{\delta}}{\alpha_{S}^{6}\beta^{4}\lambda^{2}}\right)$ in both the multiplicative and additive error. The utility guarantees using the bounds from (Høgsgaard et al., 2023) change only by logarithmic factors compared to those in Theorems 5.4 and 5.8.
@@
Differential Privacy. Differential privacy (Dwork et al., 2006b) offers privacy guarantees to individuals contributing their data to some randomized algorithm. We say that two datasets are neighboring if one can be obtained from the other by the the replacement of a single individual with another individual.
Definition 3.2 (Differential Privacy in the central model (Dwork et al., 2006b)).
@@
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\varepsilon,\delta)$-differentially private, if for all neighboring data sets $x,x^{\prime}\in\mathcal{X}$ and all outputs $S\subseteq\mathcal{Y}$ it holds that
@@
where the probabilities are over the randomness of $\mathcal{M}$. If $\delta=0$, then we simply say that the mechanism is $\varepsilon$-differentially private.
We first define the Gaussian and Laplace mechanisms, which we use later.
Lemma 3.3 (The Gaussian Mechanism (Dwork and Roth, 2014)).
@@
Lemma 3.4 (The Laplace Mechanism (Dwork et al., 2006b)).
Let $f:\mathcal{X}\to{\mathbb{R}}^{k}$ be a function and let $\varepsilon\geq 0$. The Laplace mechanism adds to each of the $k$ components of the output, noise sampled from $Lap(\Delta_{1}f/\varepsilon)$, where $\Delta_{1}f=\max_{x\sim x^{\prime}}\lvert f(x)-f(x^{\prime})\rvert$ denotes the $\ell_{1}$ sensitivity of function $f$. The Laplace mechanism is $\varepsilon$ differentially private.
Lemma 3.5 (Sequential Composition (Dwork et al., 2006b)).
A probability distribution $F$ is infinitely divisible if for any positive integer $n$, there is a random variable $S_{n}$ which can be written as $S_{n}=\sum_{i\in[n]}X_{i}$ such that each of the $X_{i}$ are random variables, independent and identically distributed, and $S_{n}$ has the probability distribution $F$.